OpenStudy (anonymous):

the positive square root of 19 is between 6&7 4 & 5 5&6 4&6 heaaalp.

jimthompson5910 (jim_thompson5910):

hints 1^2 = 1 2^2 = 4 3^2 = 9 4^2 = 16 5^2 = 25 6^2 = 36 7^2 = 49

OpenStudy (anonymous):

4 and 5

OpenStudy (anonymous):

thank you.

jimthompson5910 (jim_thompson5910):

so looking at the list of perfect squares 1, 4, 9, 16, 25, 36, 49, ... 19 is between which numbers?

OpenStudy (anonymous):

16 & 25.

jimthompson5910 (jim_thompson5910):

since 4^2 = 16 and 5^2 = 25, this means that squaring some number in between 4 and 5 will give you 19

jimthompson5910 (jim_thompson5910):

so that's why the answer is 4 and 5